OpenStudy (anonymous):

Find f(-5) if f(x) = |x + 1|. 4 5 6

OpenStudy (anonymous):

i dont understand these never learned them

OpenStudy (perl):

$$ \Large {f(x) = | x + 1 | \\f(-5) = | -5 + 1 | } $$

OpenStudy (anonymous):

would u add -5+1 then divide that answer with -5?

OpenStudy (perl):

$$\Large {f(x) = | x + 1 | \\f(-5) = | -5 + 1 |= | -4 | = 4 \\ }$$

OpenStudy (anonymous):

ohhh :o